# Generalizing Denoising to Non-Equilibrium Structures Improves Equivariant Force Fields

### v2

@@ 1 Introduction
Since a non-equilibrium structure has non-zero atomic forces and atoms are not confined to local energy minima, it can have more possible atomic positions than a structure at equilibrium. As shown in Figure 1(b), this can make denoising a non-equilibrium structure an ill-posed problem since there are many possible target structures. To address the issue, we propose force encoding and take the forces of the original non-equilibrium structures as inputs when denoising non-equilibrium structures. Intuitively, the forces constrain the atomic positions of a non-equilibrium structure. With the additional information, we are able to predict the original non-equilibrium structure satisfying the input forces instead of predicting any arbitrary structures as shown in Figure 1(d). Previous works on denoising equilibrium structures (Jiao et al., 2022; Zaidi et al., 2023; Liu et al., 2023; Feng et al., 2023b; a) end up being a special case where the forces of original structures are close to zero as in Figure 1(e).
@@ 2 Related Works
Please refer to Section A.2 for discussion on equivariant networks.
@@ 3.2.2 Force Encoding
On the other hand, it might not be that intuitive to encode forces in invariant networks since their internal latent representations such as node embeddings and edge embeddings are scalars instead of type-$L$ vectors or geometric tensors. One potential manner of encoding forces in latent representations is to project them into edge embeddings by taking inner products between forces and edge vectors of relative positions. This process is the inverse of how GemNet-OC (Gasteiger et al., 2022) decodes forces from latent representations. Since equivariant networks have been shown to outperform invariant networks on atomistic datasets and are a more natural fit to encoding forces, we mainly focus on equivariant networks and how DeNS can further advance their performance.
@@
Figure 2: Training process when incorporating DeNS as an auxiliary task. The pseudocode for training with DeNS can be found in Section E. “Equivariant GNN”, “energy head”, “force head” and “noise head” are shared across (a), (b) and (c). For each batch of structures, we use the original task (a) for some structures and DeNS ((b) or (c)) for the others. Using partially corrupted structures as in (c) is empirically better than (b). We note that the force label $F(S_{\text{non-eq}})$ and energy label $E(S_{\text{non-eq}})$ used in (b) and (c) are the same as those in (a) and that training with DeNS does not require any additional data.
@@ 3.2.3 Training with DeNS
where $\mathcal{L}_{\text{DeNS}}$ denotes the loss function of denoising as defined in Equation 4. We note that we also encode forces as discussed in Section 3.2.2 to predict the energy of $S_{\text{non-eq}}$, and we share the energy prediction head across Equation 1 and Equation 6. The loss function introduces another hyper-parameter $\lambda_{\text{DeNS}}$, DeNS coefficient, controlling the relative importance of the auxiliary task. Besides, the process of corrupting structures also results in another hyper-parameter $\sigma$ as shown in Equation 2. We provide the pseudocode in Section E. We note that the force encoding in DeNS only relies on force labels on the training set and we do not use any force labels on the validation or testing sets.
@@ 4.1 OC20 Dataset
The primary task in OC20 is Structure to Energy and Forces (S2EF), which is to predict the energy and per-atom forces given an equilibrium or non-equilibrium structure from any point in the trajectory. These predictions are evaluated on energy and force mean absolute error (MAE). Once a model is trained for S2EF, it is used to run structural relaxations from an initial structure using the predicted forces till a local energy minimum is found. The energy predictions of these relaxed structures are evaluated on the Initial Structure to Relaxed Energy (IS2RE) task.
@@ 4.2 OC22 Dataset
The Open Catalyst 2022 (OC22) dataset (Tran* et al., 2022) focuses on oxide electrocatalysis and consists of about $62$k DFT relaxations obtained with Perdew-Burke-Ernzerhof (PBE) functional calculations. One crucial difference between OC22 and OC20 is that the energy targets in OC22 are DFT total energies. DFT total energies are harder to predict but are the most general and closest to a DFT surrogate, offering the flexibility to study property prediction beyond adsorption energies. Analogous to the task definitions in OC20, the primary tasks in OC22 are S2EF-Total and IS2RE-Total. We train models on the OC22 S2EF-Total dataset, which has $8.2$M structures, and evaluate them on energy and force MAE on the S2EF-total validation and test splits. Then, we use these models to perform relaxations starting from initial structures in the IS2RE-Total test split and evaluate the predicted relaxed energies on energy MAE.
@@ Results.
{adjustwidth}
-2.5cm-2.5cm Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Training time Number of Model energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces (GPU-hours)$\downarrow$ parameters SchNet (Schütt et al., 2017) 16.0 58.5 3.5 13.4 3.5 16.9 5.6 28.6 6.9 25.2 8.7 36.9 5.2 24.7 6.1 24.3 - - DimeNet (Gasteiger et al., 2020) 8.8 21.6 3.4 8.1 2.8 10.0 4.5 16.6 5.3 9.3 5.8 16.2 4.4 9.4 5.0 13.1 - - PaiNN (Schütt et al., 2021) 6.9 14.7 - - 2.7 9.7 3.9 13.8 5.0 3.3 4.9 8.5 4.1 4.1 4.5 6.0 - - TorchMD-NET (Thölke & Fabritiis, 2022) 5.3 11.0 2.5 8.5 2.3 4.7 3.3 7.3 3.7 2.6 4.0 5.6 3.2 2.9 4.1 4.1 - - NequIP ($L_{max}=3$(Batzner et al., 2022) 5.7 8.0 - - 2.2 3.1 3.3 5.6 4.9 1.7 4.6 3.9 4.0 2.0 4.5 3.3 - - Equiformer ($L_{max}=2$) 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3 17 $3.50$M Equiformer ($L_{max}=3$) 5.3 6.6 2.5 8.1 2.2 2.9 3.2 5.4 4.4 2.0 4.3 3.9 3.7 2.1 4.3 3.4 59 $5.50$M Equiformer ($L_{max}=2$) + DeNS 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3 19 $4.00$M Equiformer ($L_{max}=3$) + DeNS 5.0 5.2 2.3 6.1 2.2 2.4 3.2 4.1 3.7 1.6 4.2 3.2 3.4 1.8 4.1 2.9 61 $6.35$M
@@
{adjustwidth}
-3.5cm-3.5cm Method Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Index Attention Layer normalization DeNS energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces 1 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3 2 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3 3 5.2 7.7 2.4 6.2 2.3 3.9 3.3 6.2 3.8 2.2 4.1 4.7 3.3 2.4 4.2 4.4 4 5.2 6.1 2.4 6.1 2.2 2.9 3.2 5.1 3.7 1.7 4.2 3.9 3.4 2.0 4.2 3.4 5 5.3 9.3 2.4 9.2 2.3 4.5 3.4 8.2 3.7 2.4 4.2 5.5 3.3 2.9 4.2 4.8 6 5.2 7.3 2.4 8.1 2.2 3.4 3.4 6.7 3.7 1.9 4.2 4.4 3.4 2.2 4.2 3.8
@@ 4.3 MD17 Dataset
The MD17 dataset (Chmiela et al., 2017; Schütt et al., 2017; Chmiela et al., 2018) consists of molecular dynamics simulations of small organic molecules. The task is to predict the energy and forces of these non-equilibrium molecules. We use $950$ and $50$ different configurations for training and validation sets and the rest for the testing set.
@@
We train Equiformer ($L_{max}=2$(Liao & Smidt, 2023) and Equiformer ($L_{max}=3$) with DeNS based on their official implementation, where $L_{max}$ denotes the maximum degree of equivariant representations. As shown in Table 4, DeNS improves the results on all molecules. Along with the results on OC20 and OC22 datasets, DeNS can generally improve the performance on force predictions with both direct and gradient methods. Particularly, Equiformer ($L_{max}=2$) trained with DeNS acheives better results on all the tasks and requires 3.1×3.1$\times$3.1 × less training time than Equiformer ($L_{max}=3$) trained without DeNS. This demonstrates that for this small dataset, training an auxiliary task and using data augmentation are more efficient and result in larger performance gain than increasing $L_{max}$ from $2$ to $3$. Additionally, we find that the gains from training DeNS as an auxiliary task are comparable to pre-training. For example, Zaidi et al. (2023) uses TorchMD-NET (Thölke & Fabritiis, 2022) pre-trained on the PCQM4Mv2 dataset and reports results on Aspirin. Their improvement in force MAE is about $17.2\%$ (Table 3 in Zaidi et al. (2023)). Training Equiformer ($L_{max}=2$) with DeNS results in $20.8\%$ improvement in force MAE without relying on another dataset. Note that we only increase training time by $10.5\%$ while their method takes much more time since PCQM4Mv2 dataset is more than 3000×3000$\times$3000 × larger than the training set of MD17. Moreover, training with DeNS enables consistent improvement in all molecules when increasing $L_{max}$ from $2$ to $3$, and Equiformer ($L_{max}=3$) trained with DeNS achieves overall best results. In contrast, when DeNS is not used, increasing $L_{max}$ from $2$ to $3$ can lead to overfitting and worse results on some molecules (i.e., Benzene and Uracil).
@@
{adjustwidth}
-3.5cm-3.5cm Method Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Force $\lambda_{E}\neq 0$ Partially corrupted Index DeNS encoding in Eq. 6 structures energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces 1 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3 2 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3 3 8.6 9.1 2.3 6.3 2.3 3.3 3.2 5.8 7.7 6.1 5.2 10.6 3.7 2.0 5.5 6.5 4 5.1 5.8 2.3 6.1 2.3 2.6 3.2 4.5 3.7 1.7 4.8 3.6 3.7 1.9 4.2 3.3 5 5.3 6.9 2.4 6.3 2.3 3.2 3.3 5.4 3.7 2.0 4.8 4.2 4.0 2.0 4.2 3.8
@@ 5 Conclusion
In this paper, we propose to use denoising non-equilibrium structures (DeNS) as an auxiliary task to better leverage training data and improve performance on original tasks of energy and force predictions. Denoising non-equilibrium structures can be an ill-posed problem since there are many possible target structures. To address the issue, we propose force encoding and take the forces of original structures as inputs to specify which non-equilibrium structures we are denoising. With force encoding, DeNS successfully improves the performance on original tasks when it is used as an auxiliary task. We conduct extensive experiments on OC20, OC22 and MD17 datasets to demonstrate that DeNS can boost the performance on energy and force predictions across datasets of various scales with minimal increase in training cost and is applicable to many equivariant networks.
@@
  • E Pseudocode for training with DeNS
  • F Visualization of corrupted structures

### v3

@@ 1 Introduction
Since a non-equilibrium structure has non-zero atomic forces and atoms are not confined to local energy minima, it can have more possible atomic positions than a structure at equilibrium. As shown in Figure 1(b), this can make denoising a non-equilibrium structure an ill-posed problem since there are many possible target structures. To address the issue, we propose force encoding and take the forces of the original non-equilibrium structures as inputs when denoising non-equilibrium structures. Intuitively, the forces constrain the atomic positions of a non-equilibrium structure. With the additional information, we are able to predict the original non-equilibrium structure satisfying the input forces instead of predicting any arbitrary structures as shown in Figure 1(d). Previous works on denoising equilibrium structures (Godwin et al., 2022; Jiao et al., 2022; Zaidi et al., 2023; Liu et al., 2023; Feng et al., 2023b; a) end up being a special case where the forces of original structures are close to zero as in Figure 1(e).
@@ 2 Related Works
Please refer to Section A.2 for discussion on equivariant networks. In addition, since we mainly focus on applying the proposed DeNS to Equiformer series (Liao & Smidt, 2023; Liao et al., 2023), we provide a brief introduction to them in Section A.3.
@@ 3.2.2 Force Encoding
On the other hand, it might not be that intuitive to encode forces in invariant networks since their internal latent representations such as node embeddings and edge embeddings are scalars instead of type-$L$ vectors or geometric tensors. One potential manner of encoding forces in latent representations is to project them into edge embeddings by taking inner products between forces and edge vectors of relative positions. This process is the inverse of how GemNet-OC (Gasteiger et al., 2022) decodes forces from latent representations. Since equivariant networks have been shown to outperform invariant networks on atomistic datasets and are a more natural fit to encoding forces, we mainly focus on equivariant networks and how DeNS can further advance their performance. As for other work (Duval et al., 2023) leveraging frame averaging to achieve equivariance through data transformations, we can follow how unit cell Cartesian coordinates are projected in their framework to encode forces when optimizing for DeNS. Specifically, we compute one set of frame axes using only 3D atomic positions and project the input forces to the frame axes. The projected forces remain the same under any $E(3)$ transformation and enable using unconstrained functions to encode the input forces.
@@
Figure 2: Training process when incorporating DeNS as an auxiliary task. The pseudocode for training with DeNS can be found in Section F. “Equivariant GNN”, “energy head”, “force head” and “noise head” are shared across (a), (b) and (c). For each batch of structures, we use the original task (a) for some structures and DeNS ((b) or (c)) for the others. Using partially corrupted structures as in (c) is empirically better than (b). We note that the force label $F(S_{\text{non-eq}})$ and energy label $E(S_{\text{non-eq}})$ used in (b) and (c) are the same as those in (a) and that training with DeNS does not require any additional data.
@@ 3.2.3 Training with DeNS
where $\mathcal{L}_{\text{DeNS}}$ denotes the loss function of denoising as defined in Equation 4. We note that we also encode forces as discussed in Section 3.2.2 to predict the energy of $S_{\text{non-eq}}$, and we share the energy prediction head across Equation 1 and Equation 6. The loss function introduces another hyper-parameter $\lambda_{\text{DeNS}}$, DeNS coefficient, controlling the relative importance of the auxiliary task. Besides, the process of corrupting structures also results in another hyper-parameter $\sigma$ as shown in Equation 2. We provide the pseudocode in Section F. We note that the force encoding in DeNS only relies on force labels on the training set and we do not use any force labels on the validation or testing sets.
@@ 4.1 OC20 Dataset
The primary task in OC20 is Structure to Energy and Forces (S2EF), which is to predict the energy and per-atom forces given an equilibrium or non-equilibrium structure from any point in the trajectory. These predictions are evaluated on energy and force mean absolute error (MAE). Most of the previous works use direct methods to predict forces on OC20, and we follow this practice and investigate how DeNS can improve direct methods for force prediction. Once a model is trained for S2EF, it is used to run structural relaxations from an initial structure using the predicted forces till a local energy minimum is found. The energy predictions of these relaxed structures are evaluated on the Initial Structure to Relaxed Energy (IS2RE) task.
@@ 4.2 OC22 Dataset
The Open Catalyst 2022 (OC22) dataset (Tran* et al., 2022) focuses on oxide electrocatalysis and consists of about $62$k DFT relaxations obtained with Perdew-Burke-Ernzerhof (PBE) functional calculations. One crucial difference between OC22 and OC20 is that the energy targets in OC22 are DFT total energies. DFT total energies are harder to predict but are the most general and closest to a DFT surrogate, offering the flexibility to study property prediction beyond adsorption energies. Analogous to the task definitions in OC20, the primary tasks in OC22 are S2EF-Total and IS2RE-Total. We train models on the OC22 S2EF-Total dataset, which has $8.2$M structures, and evaluate them on energy and force MAE on the S2EF-total validation and test splits. Same as OC20, we use direct methods to predict forces here. Then, we use these models to perform relaxations starting from initial structures in the IS2RE-Total test split and evaluate the predicted relaxed energies on energy MAE.
@@ Results.
Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Training time Number of
Model energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces (GPU-hours)$\downarrow$ parameters
SchNet (Schütt et al., 2017) 16.0 58.5 3.5 13.4 3.5 16.9 5.6 28.6 6.9 25.2 8.7 36.9 5.2 24.7 6.1 24.3 - -
DimeNet (Gasteiger et al., 2020) 8.8 21.6 3.4 8.1 2.8 10.0 4.5 16.6 5.3 9.3 5.8 16.2 4.4 9.4 5.0 13.1 - -
PaiNN (Schütt et al., 2021) 6.9 14.7 - - 2.7 9.7 3.9 13.8 5.0 3.3 4.9 8.5 4.1 4.1 4.5 6.0 - -
TorchMD-NET (Thölke & Fabritiis, 2022) 5.3 11.0 2.5 8.5 2.3 4.7 3.3 7.3 3.7 2.6 4.0 5.6 3.2 2.9 4.1 4.1 - -
NequIP ($L_{max}=3$(Batzner et al., 2022) 5.7 8.0 - - 2.2 3.1 3.3 5.6 4.9 1.7 4.6 3.9 4.0 2.0 4.5 3.3 - -
Equiformer ($L_{max}=2$) 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3 17 $3.50$M
Equiformer ($L_{max}=3$) 5.3 6.6 2.5 8.1 2.2 2.9 3.2 5.4 4.4 2.0 4.3 3.9 3.7 2.1 4.3 3.4 59 $5.50$M
Equiformer ($L_{max}=2$) + DeNS 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3 19 $4.00$M
Equiformer ($L_{max}=3$) + DeNS 5.0 5.2 2.3 6.1 2.2 2.4 3.2 4.1 3.7 1.6 4.2 3.2 3.4 1.8 4.1 2.9 61 $6.35$M
@@
Method Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil
Index Attention Layer normalization DeNS energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces
1 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3
2 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3
3 5.2 7.7 2.4 6.2 2.3 3.9 3.3 6.2 3.8 2.2 4.1 4.7 3.3 2.4 4.2 4.4
4 5.2 6.1 2.4 6.1 2.2 2.9 3.2 5.1 3.7 1.7 4.2 3.9 3.4 2.0 4.2 3.4
5 5.3 9.3 2.4 9.2 2.3 4.5 3.4 8.2 3.7 2.4 4.2 5.5 3.3 2.9 4.2 4.8
6 5.2 7.3 2.4 8.1 2.2 3.4 3.4 6.7 3.7 1.9 4.2 4.4 3.4 2.2 4.2 3.8
@@
Method Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil
Force $\lambda_{E}\neq 0$ Partially corrupted
Index DeNS encoding in Eq. 6 structures energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces
1 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3
2 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3
3 8.6 9.1 2.3 6.3 2.3 3.3 3.2 5.8 7.7 6.1 5.2 10.6 3.7 2.0 5.5 6.5
4 5.1 5.8 2.3 6.1 2.3 2.6 3.2 4.5 3.7 1.7 4.8 3.6 3.7 1.9 4.2 3.3
5 5.3 6.9 2.4 6.3 2.3 3.2 3.3 5.4 3.7 2.0 4.8 4.2 4.0 2.0 4.2 3.8
@@ 4.3 MD17 Dataset
The MD17 dataset (Chmiela et al., 2017; Schütt et al., 2017; Chmiela et al., 2018) consists of molecular dynamics simulations of small organic molecules. The task is to predict the energy and forces of these non-equilibrium molecules. Following previous works, we adopt gradient methods for force prediction. We use $950$ and $50$ different configurations for training and validation sets and the rest for the testing set.
@@
We train Equiformer ($L_{max}=2$(Liao & Smidt, 2023) and Equiformer ($L_{max}=3$) with DeNS based on their official implementation, where $L_{max}$ denotes the maximum degree of equivariant representations. As shown in Table 4, DeNS improves the results on all molecules. Along with the results on OC20 and OC22 datasets, DeNS can generally improve the performance on force predictions with both direct (i.e., OC20 and OC22) and gradient (i.e., MD17) methods. Particularly, Equiformer ($L_{max}=2$) trained with DeNS acheives better results on all the tasks and requires 3.1×3.1$\times$3.1 × less training time than Equiformer ($L_{max}=3$) trained without DeNS. This demonstrates that for this small dataset, training an auxiliary task and using data augmentation are more efficient and result in larger performance gain than increasing $L_{max}$ from $2$ to $3$. Additionally, we find that the gains from training DeNS as an auxiliary task are comparable to pre-training. For example, Zaidi et al. (2023) uses TorchMD-NET (Thölke & Fabritiis, 2022) pre-trained on the PCQM4Mv2 dataset and reports results on Aspirin. Their improvement in force MAE is about $17.2\%$ (Table 3 in Zaidi et al. (2023)). Training Equiformer ($L_{max}=2$) with DeNS results in $20.8\%$ improvement in force MAE without relying on another dataset. Note that we only increase training time by $10.5\%$ while their method takes much more time since PCQM4Mv2 dataset is more than 3000×3000$\times$3000 × larger than the training set of MD17. Moreover, training with DeNS enables consistent improvement in all molecules when increasing $L_{max}$ from $2$ to $3$, and Equiformer ($L_{max}=3$) trained with DeNS achieves overall best results. In contrast, when DeNS is not used, increasing $L_{max}$ from $2$ to $3$ can lead to overfitting and worse results on some molecules (i.e., Benzene and Uracil). Besides, we report simulation-based results in Section D.3.
@@ 5 Conclusion
In this paper, we propose to use denoising non-equilibrium structures (DeNS) as an auxiliary task to better leverage training data and improve performance on original tasks of energy and force predictions. Denoising non-equilibrium structures can be an ill-posed problem since there are many possible target structures. To address the issue, we propose force encoding and take the forces of original structures as inputs to specify which non-equilibrium structures we are denoising. With force encoding, DeNS successfully improves the performance on original tasks when it is used as an auxiliary task. We conduct extensive experiments on OC20, OC22 and MD17 datasets to demonstrate that DeNS can boost the performance on energy and force predictions across datasets of various scales with minimal increase in training cost and is applicable to many equivariant networks. Finally, we note that the proposed DeNS is general and can be directly applied to other atomistic datasets containing non-equilibrium structures. Take Materials Project (Jain et al., 2013) for example. Similar to OC20 and OC22 datasets, for each entry in the Materials Project database, the Materials Project Trajectory (MPtrj) dataset (Deng et al., 2023) contains the corresponding relaxation trajectory between initial and relaxed structures. Most structures in MPtrj are non-equilibrium and have energy and force labels. Therefore, we can apply DeNS to MPtrj dataset in the same way as OC20 and OC22 datasets.
Acknowledgement
We thank Larry Zitnick, Xinlei Chen, Zachary Ulissi, Saro Passaro, Anuroop Sriram, and Brandon Wood for helpful discussions. We acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center (Reuther et al., 2018) for providing high performance computing and consultation resources that have contributed to the research results reported within this paper.
Yi-Lun Liao and Tess Smidt were supported by DOE ICDI grant DE-SC0022215.
@@ Acknowledgement
    3. D.3 Additional simulation-based results
  • E Pseudocode for force encoding
  • F Pseudocode for training with DeNS
  • G Visualization of corrupted structures